Ajit has a certain average for 9 innings. In the 10th inning he scores 100 sun`s thereby increasing his average by 8 runs. His new average is.??

Grade:12th pass

1 Answers

sahil
6 years ago
Sol:Let ajit score in 9 matches be X.Now his avg =x/9,He score 100 in tenth match so now avg =(X+100)/10;and difference of averages is 8 so A/Q :8=(100+X)/10-X/9 OR 8=(900+9X-10X)/90 OR X= 180 .So new avg was (100+X)/10 putting x value (280)/10 or 28 . Hence required new avg is 28.Hope it helps if yes please give approval